if f(x)=3x+6 and g(x)=2x^2-x+3 what is g[f(2)]?
Answer by stanbon(75887)

if f(x)=3x+6 and g(x)=2x^2-x+3 what is g[f(2)]
----
g[f(2)] = g[3*2+6] = g[12] = 2*12^2-12+3 = 279
====================
Cheers,
Stan H.